# Missing Pieces: How Framing Uncertainty Impacts Longitudinal Trust in AI Decision Aids — A Gig Driver Case Study

### v1

@@ 4.2. Tool Conditions
Reference condition. Participants were only shown their mean weekly estimated earnings and their mean estimated earnings for each hour during the week.
@@
Daily estimates. To assess the effect of increasing granularity but not error rate visibility, participants were shown their mean estimated earning for each day instead of their mean estimated weekly earning. The schedule still showed mean estimated earnings for each hour during the week.
Ranged estimates. To assess the effect of increasing error rate visibility, participants were shown mean, pessimistic, and optimistic estimated earnings for the entire week, as well as for each hour during the week.
@@
Ranged and hedged estimates. To assess the effect of hedging on error rate visibility, the textual description of the estimates was changed from Condition 3. Instead of “Based on historical data, it is estimated that you will earn”, Condition 4 states “On average, based on historical data, a driver following this schedule will earn”.
5. Pilot Interviews
To assess the utility of the AI gig recommendation tool design introduced in Section 3, we created a prototype based on them. We then conducted a series of interviews to understand gig drivers’ needs and how well the prototype aligned with them. The interview methodology was approved by our Institutional Review Board (IRB).
5.1. Tool Design
The constraint page’s design (Figure S6) was largely similar to the final design. However, the prototype had a monotone colour scheme and had instructions that were worded less clearly, which were improved based on feedback from the pilot (Section 5.3). Additionally, due to technical limitations, dropdowns were used in place of text boxes, and no input validation was present. However, we consider the impact of these limitations to be minor given that the interviewer controlled the page.
The schedule page’s design (Figure S7) was kept static for the pilot to gather more uniform feedback from participants. Like the constraint page, the prototype differed from the final design in its colour scheme and clarity of wording. The prototype was most similar to Condition 3 (Section 4.2) in that it displayed a range consisting of mean (“On an average week…”), pessimistic (“On a bad week…”), and optimistic (“On a good week…”) weekly earnings above the tabular schedule. Unlike the final design of Condition 3, however, ranges were not shown for hourly estimates; this was added based on feedback from the pilot (Section 5.3).
5.2. Methodology
Participants began by completing a web-based consent form and a demographics survey that collected their age, gender, and education level. This form was separate from the one for the longitudinal trials (Section 6.1.1). After completing the web-based portion, participants were invited via email to complete 20–30-minute audio-recorded Zoom interviews conducted by a single author. Participants were compensated with a $10 Amazon gift card.
In the first 5–10 minutes, the interview focused on formatively understanding drivers’ needs and motivations. We asked the participants a series of questions about their needs and motivations, including what they liked or disliked about driving for gig platforms, their typical goals and routines when driving, and the extent to which they would find schedule recommendations useful, along with demographic questions.
In the last 15–20 minutes, the interview focused on evaluatively understanding how well the tool met drivers’ needs. To ensure a consistent experience, the Figma prototype was opened in the interviewer’s browser and shown to participants in a screensharing session. First, on the constraint page, the participant was asked to work with the interviewer to interact with the page, entering the constraints as if they were using the tool for their actual planning. Then, on the schedule page, the participant was shown a schedule with mocked earning estimates. Finally, the participant was asked about their overall opinions of the tool. The full interview script is shown in Section B.1.
Transcripts for the interviews were automatically generated by Zoom. The author who conducted the interviews first reviewed and corrected these transcripts, then performed structural coding (Saldaña, 2015a). Afterwards, two authors separately used QualCoder 3.3 (Curtain, 2023) to perform open coding (through descriptive, process, and in vivo coding (Saldaña, 2015a)) and axial coding (Saldaña, 2015b). The two authors then met to reconcile their codes and construct a unified codebook. Finally, the interviewing author re-applied the updated set of codes. Following (Cai et al., 2019b), we did not compute inter-rater reliability, since our grounded analysis used codes as an intermediary to generate theory.
5.2.1. Participants
As with the longitudinal trials (Section 6.1.3), participants were recruited from the user base of Gridwise, a mobile assistant app for gig drivers, in July 2023. We chose this user base instead of recruiting from a particular platform (Zhang et al., 2023) or from social media (Zhang et al., 2022), so as to capture the needs of a diverse set of drivers. Gridwise distributed recruitment messages to 500 users, but otherwise did not interact with participants. Recipients were sampled from users in the United States who had completed at least one gig in a platform linked to the Gridwise app over the week preceding recruitment. We recruited 4 interview participants, hereafter designated as P1–P4:
5.3. Results
The four participants reported a diversity of motivations and routines. While all four participants had specific earning goals, P1, P2, and P4 considered their goals to be important and valued the time flexibility of gig work, whereas P3 was more motivated by the opportunity for human interaction. P1, P2 and P3 had typical times that they drive at, but P1, P2, and P4 also adjusted their driving schedules based on gig availability. All four participants had encountered difficulties in planning due to the unpredictability of rider demand and/or driver supply (with P1, P2, and P3 feeling that gig platforms provide insufficient information), and indicated that they would find schedule recommendations to be useful.
All four participants found the initial design of the tool to be generally understandable, and felt that it would be generally useful for drivers in planning. P1 and P2 liked the fact that the tool presents information to them in a way that reduces the need for guesswork while driving. In particular, P1 suggested that the tool would help mitigate a catch-22: it is not possible to view gig demand information in DoorDash without exiting their Dash (scheduled work period), but doing so seemingly deprioritises them.
On the constraint page, P1 and P4 indicated that the questions aligned well with their goals. P1 and P3 suggested that they would not set the constraints to perfectly align with their routines, so as to receive more information from the tool. Meanwhile, on the schedule page, all four participants liked the estimated hourly earnings, with P1, P2, and P3 indicating that they would be helpful in deciding whether or not to work at particular times of day. Yet, P2, P3, and P4 acknowledged that the estimates were only guesses. P1 and P4 also liked the range of weekly earnings, but P3 felt it assumed they would follow the recommended schedule perfectly. P2 and P3 mentioned that ranges for hourly earnings would be useful.
More granular constraints. P1, P3, and P4 all suggested ways to limit the scope of the historical gigs used to estimate their earnings. P3 wanted the tool to clarify that the historical data was limited to the city they drive in, and also how recent the data was. P1 and P3 wanted to limit the maximum distance of the historical gigs from their starting point. P4, who works for less popular delivery platforms, wanted to select which platforms the historical gigs came from, and indicated that this would improve their control.
Feedback on performance. P1 and P2 both wanted a way to compare the tool’s estimates with their actual earnings. Regardless of how the estimates compared to reality, P1 suggested that they would be motivating; both P1 and P2 remarked that they have gamified their gig-driving experiences to compare against either themselves or others. P2 also wanted to compare the estimates with their expenses, and P3 wanted a way to view the overall supply of drivers.
6. Longitudinal Trials
After our pilot study, through which we verified the problem formulation and improved the design of our AI decision aid (Section 5), we sought to deploy the decision aid to a larger sample of users. Our goal was to measure how their trust in the tool would evolve over time and vary with different framings of uncertainty. To this end, we conducted a longitudinal trial in which participants interacted with the tool for 7 days over a 14-day period, with the longer time window meant to accommodate variability in participants’ availability. The methodology for this longitudinal trial was approved by our Institutional Review Board (IRB).
6.1. Methodology
6.1.1. Trial Design
The trial was administered through a custom website built using Django 4.1 (Django, 2023), which was designed to be functional on both computers and mobile devices. Crucially, the website was designed to be visually generic and distinct from any apps or websites that participants may use to access gig platforms; this was done to avoid the possibility that participants’ experiences with any individual platform could have biased their perception of the tool. Prospective participants received a link to the website from the recruitment message, which led them to complete the consent form and begin the trial.
On Day 0 of the trial, participants first completed the 12-question Survey 1 — Formative Questions (LABEL:{sec:app:survey:1}). These questions were based on our formative interviews about participants’ needs and motivations from the pilot interviews (Section 5.2). To filter the data used to generate estimates, we also asked participants about the region they drive in and the platforms they work for.
Next, they were directed to a page on which they interacted with the schedule recommendation tool. To mitigate response bias (Dell et al., 2012; Ko et al., 2015), we avoided referring to the tool as having been designed by us; isolated the tool in an iframe; and differentiated the fonts of the website and tool. First, participants filled in their availability and goals on the constraint page. Second, on the schedule page, we assigned participants to one of the 4 conditions to view their schedule. Assignment was done such that each condition had an approximately equal number of participants.
The estimates were generated using gig data from August 2023 in four cities: Los Angeles, New York, Chicago, and Houston. For each city, our data included approximately 100 000–300 000 samples distributed evenly across times and weekdays. Earnings for each time slot (weekday , hour ) were estimated by the mean of what drivers historically earned in this slot, filtered to the participant’s city and platforms. We used this simple, static estimator to focus on the effects of framing uncertainty in the estimates. Our results are intended, however, to generalise to any possible estimator for , including those that generate dynamic estimates based on machine learning. To this end, we intentionally chose not to emphasise to the participants that the schedule page was generated using AI or optimisation, and did not provide any additional information about the data used to generate the estimates.
We sought to measure participants’ trust in the tool throughout the longitudinal trial. As reviewed by (Kohn et al., 2021), internal trust is typically measured by self-report measures. Two widely-used instruments for self-reported trust are (Madsen and Gregor, 2000)’s Human-Computer Trust Questionnaire and (Jian et al., 2000)’s Trust in Automation Scale. The former is a 25-item scale that measures 5 facets of trust (Perceived Understandability, Perceived Technical Competence, Perceived Reliability, Personal Attachment, and Faith) using 5 questions each. Meanwhile, the latter is a 12-item scale that measures both trust and distrust. Following the recommendations of (Kohn et al., 2021), we adapted these validated instruments to our specific context instead of creating custom measures of trust (see comments in Appendix A). To reduce user burden, we considered it infeasible to repeatedly administer these instruments in their entirety, and instead extracted subsets of them for our surveys as recommended by (Kohn et al., 2021). The first measurement of trust took place immediately after participants saw the recommended schedule on Day 0. Survey 2 — Pre-Interaction Trust (Section A.2) consisted of 5 items, one taken from each of (Madsen and Gregor, 2000)’s 5 facets of trust.
After completing Survey 2, participants began their 7 days of interaction with the tool starting on their next available date. On each of Day 1 to Day 7, participants first received their recommended schedule for that particular day, sent via an email scheduled for 30 minutes before the start of their indicated availability. This was intended to present participants with information from the tool right as they were deciding their driving schedule. During the day, participants continued uninterrupted with their driving activity; we repeatedly emphasised that compliance with the recommended schedule was not a condition of full participation. Finally, at the end of their indicated availability, a second scheduled email sent them a link to Survey 3 — Daily Trust (Section A.3). This was an 8-question survey that included items extracted from 3 of (Madsen and Gregor, 2000)’s 5 facets of trust, along with questions on how the participant’s earnings compared to their expectations and the tool’s estimates, and to what extent they intended to rely on the tool for the following day.
If the participant indicated that they intended to continue relying on the tool, they were then presented with a daily variant of the schedule page, showing the subset of the recommended schedule corresponding to the following day. We also allowed them to alter their availability for the coming day using a daily variant of the constraint page. Updated schedules were generated by fixing the recommended time slots for previous days using equality constraints and then re-solving Equation 1 or Equation 2. On the other hand, if the participant’s answer indicated that they intended to pause their interaction for one day, a scheduled email was sent on the following day, which prompted them to either review the next day’s schedule or to pause for an additional day.
On Day 7, we removed the last question of Survey 3 and added Survey 4 — Post-Interaction Trust (Section A.4), a series of 10 questions intended as a retrospective measurement of trust after participants had completed interaction with the tool. This survey included 5 questions drawn from (Jian et al., 2000)’s instrument. We chose not to include questions from this instrument earlier in the trial, as we felt that questions measuring distrust could have biased participants’ reliance. After completing Survey 4, participants were sent a final email that invited them to participate in the interview.
@@ 6.1.1. Trial Design
Based on a pilot of the longitudinal trial conducted with 7 participants in August 2023, we determined that the sequence of surveys on Day 0 took an average of 14 minutes and 27 seconds (including idle time), Survey 3 took an average of 2 minutes and 23 seconds per day, and Survey 4 took an average of 2 minutes and 21 seconds. The results from this pilot were excluded from analysis, and the 7 participants were excluded from future recruitment. These figures led us to set the compensation as an $40 Amazon gift card for full study completion: $6 for the Day 0 surveys, $2 for each daily survey, and a $20 completion bonus. For this payment structure, we drew from (van Berkel and Kostakos, 2021)’s recommendation of micro-compensation. However, we made one payment upon either study completion or the passage of 14 days to mitigate effects on ecological validity.
6.1.2. Interview Design
For participants who indicated their desire to be interviewed, an audio-recorded Zoom interview of 20–30 minutes was conducted by a single author. The interview was not a condition of full study completion. Participants were compensated separately with a $20 Amazon gift card.
Overall, the interview focused on assessing dimensions of participants’ experiences that were not evident from the surveys, such as the thought processes underlying their perceptions of the tool and their decisions regarding reliance and compliance. The interview began with questions about participants’ motivations and routines, which led into questions assessing the constraint page’s alignment with their decision-making process as well as their perceived control over the schedule through this page. Next, participants were asked about their understanding of the schedules, the extent of their reliance on and compliance with them, their perceptions of the outcomes, and whether these changed throughout their 7 days of interaction with the tool. Another set of questions then focused on the estimates, including perceptions of how accurate they were and whether participants would have preferred a different condition. We also asked participants to recall a specific day of interaction with the tool, in terms of how the schedule and estimates affected their behaviour for that day and for the following day. Finally, participants were asked for their overall thoughts on the tool.
6.1.3. Participants
@@
Participants were recruited from the user base of Gridwise in September 2023. We chose to recruit from this user base again to access a larger and more diverse sample of both historical data and participants. Gridwise distributed recruitment messages to users who had completed at least one gig in DoorDash, Grubhub, Instacart, Lyft, Uber, or Uber Eats (the six platforms for which historical data was available) over the month preceding recruitment, and resided in one of the the four cities with the historical data used to generate the tool’s earning estimates (see Section 6.1.1): Los Angeles, New York, Chicago, and Houston.
Bar graph showing the number of participants from Day 0 to Day 7, followed by the Interview. Day 0 had Survey 1, the constraint and schedule pages, and Survey 2 as associated activities. On Day 0, the four conditions respectively had 13, 14, 12, and 12 participants. Days 1–7 each had the daily schedule and Survey 3 as associated activities; Day 7 also had Survey 4. Over the following week, retention for Conditions 2 and 4 remained high, but retention dropped substantially for Conditions 1 and 3. By Day 7, the four conditions respectively had 6, 10, 6, and 11 participants. The Interview had 0, 3, 1, and 3 participants for the four conditions.
@@
Of the recipients, 51 completed at least the sequence of surveys on Day 0 and thus qualified for compensation. Among them, 25 (49%) were from Los Angeles, 10 (19.6%) were from New York, 8 (15.7%) were from Chicago, and 8 (15.7%) were from Houston; 4 (7.8%) were aged 18–24, 15 (29.4%) were aged 25–34, 22 (43.13%) were aged 35–44, 8 (15.7%) were aged 45–54, and 2 (3.9%) were aged over 55; 34 (66.7%) were male, 15 (29.4%) were female, and 1 (2%) was non-binary; 5 (9.8%) had a graduate degree, another 16 (31.4%) had an undergraduate degree, another 11 (21.6%) had a professional degree, and another 19 (37.3%) had a high school degree. We did not collect race and ethnicity data, as this is a rare practice that is not directly relevant to our research questions (Chen et al., 2023).
Among these, 7 of these completed the interview and received an additional $20 of compensation. They were:

### v2

@@ 3.3. Interface Conditions
Base condition. Users were only shown their mean weekly estimated earnings, and their mean estimated earnings for each hour during the week.
@@
Daily estimates. To assess the effect of introducing additional information irrelevant to uncertainty, users were shown their mean estimated earning for each day instead of their mean estimated weekly earning. As in (B), the schedule still showed mean estimated earnings for each hour.
Ranged estimates. To assess the effect of exposing uncertainty through range-based estimates (similar to (Prabhudesai et al., 2023)), users were shown mean, pessimistic, and optimistic estimates for hourly and weekly earnings.
@@
Ranged and hedged estimates. To assess the effect of exposing uncertainty through lexical hedging (similar to (Kim et al., 2024; Zhang et al., 2022b)), the textual description of the estimates was changed from (R). Instead of “Based on historical data, it is estimated that you will earn”, (RH) states “On average, based on historical data, a driver following this schedule will earn”.
4. Longitudinal User Study Design
To address Research Question 1, we conducted a longitudinal, in situ user study in which gig drivers repeatedly interacted with our schedule recommendation tool, and we measured their trust and reliance over these interactions. Our participants used the tool for 7 days over a 14-day period, with the longer time window meant to accommodate variability in participants’ availability. The methodology for this study was approved by our Institutional Review Board (IRB). Figure 2 illustrates the flow of the user study; we detail each day’s study activities in Section 4.2.
@@ 4. Longitudinal User Study Design
Based on a pilot conducted with 7 participants in August 2023, we determined that the Intake Survey, Pre-Survey, and tool interaction on Day 0 took an average of 14 minutes and 27 seconds, the End-of-Day Survey took an average of 2 minutes and 23 seconds per day, and the Post-Survey took an average of 2 minutes and 21 seconds. Based on van Berkel and Kostakos (2021)’s recommendation of micro-compensation, this led us to set the compensation as an Amazon gift card with $6 for the Day 0 surveys, $2 for each daily survey, and a $20 completion bonus ($40 for full study completion). We made one payment upon study completion or the passage of 14 days.
4.1. Participants and Data Sources
@@
Participants were recruited from the user base of Gridwise, a mobile assistant app for gig drivers, in September 2023. We chose to recruit from this user base to access a relatively large and diverse sample of both historical data and participants. Gridwise distributed recruitment messages to users who (1) had completed at least one gig in DoorDash, Grubhub, Instacart, Lyft, Uber, or Uber Eats over the month preceding recruitment, and (2) resided in one of the four cities with the historical data used to generate the tool’s earnings estimates: Los Angeles, New York, Chicago, and Houston. These were the platforms and cities for which historical data was available.
Accordingly, we generated estimates using gig data from August 2023 in each of these four cities. For each city, our data included approximately 100 000–300 000 gig records distributed evenly across times and weekdays. Hourly earnings were estimated by the mean of what drivers historically earned in this slot, filtered to the participant’s city and platforms. We used this static estimator to focus on the effects of exposing uncertainty in the estimates. Thus, we did not emphasise to participants that the schedule page was generated using AI or optimisation, and did not provide any additional information about the data used to generate the estimates.
4.2. User Study Activities
Day 0: Pre-Interaction
Participants received a link to the study website from an email-distributed recruitment message. After the consent form, they completed the first of four surveys, the Intake Survey (Section A.1). This 12-question survey asked about their needs and motivations as gig drivers, along with demographics.
Next, participants were directed to interact with the tool, which we displayed in an iframe to mitigate response bias (Dell et al., 2012; Ko et al., 2015). They entered their constraints on the constraint page, and received the tool’s recommended schedule for the entire week on the schedule page. Participants were randomly and uniformly assigned to one of the four conditions for the schedule page (Section 3.3), such that each condition had an approximately equal number of participants.
Lastly, participants completed the second of four surveys, the Pre-Survey (Section A.2), which was a 5-question survey measuring trust before interaction with the tool (Section 5.1).
Days 1–7: Interaction
Next, participants began their 7 days of interaction with the tool, beginning on the next day of the week for which they indicated they were available to drive.
On each day, participants first received their recommended schedule for that day, sent via an email scheduled for 30 minutes before the start of their indicated availability. Thus, the tool’s outputs were displayed right as they were deciding their driving schedules. During the day, participants independently made decisions about their driving activity; we emphasised that compliance with the recommended schedule was not a condition of full participation.
At the end of each participant’s indicated availability, a second scheduled email sent them a link to the End-of-Day Survey (Section A.3). This was an 8-question survey that measured their trust in the tool for that day, and their intention to rely on the tool for the next day (Section 5.1). If the participant intended to continue relying on the tool, they were then presented with a daily variant of the schedule page. Here, they could review the recommended schedule for the following day, and revise their constraints for the day as desired. Updated schedules were generated by fixing the recommended time slots for previous days using equality constraints and then re-solving the optimisation problem. However, if the participant intended to pause their interaction for one day, an email was sent on the next day, which prompted them to either review the next day’s schedule or to pause for an additional day.
Day 7: Post-Interaction
On the final day, we removed the last question of the End-of-Day Survey and added the Post-Survey (Section A.4). This was a 10-question survey that retrospectively measured participants’ trust and distrust in the tool over the course of the entire user study (Section 5.1). After completing the Post-Survey, participants were sent a final email that invited them to participate in an optional Exit Interview (Section 4.3).
4.3. Interview Procedure
For participants who indicated their desire to be interviewed, an audio-recorded Zoom interview of 20–30 minutes was conducted by a single author. The interview focused on assessing dimensions of participants’ experiences that were not evident from the surveys. We began with questions about participants’ motivations and routines, which led into questions assessing the constraint page’s alignment with their decision-making process. Next, we asked participants about the schedule page, including how the recommended schedules factored into their decision-making and how it impacted the outcomes of their driving. Further questions focused on the earnings estimates, including perceptions of their accuracy and whether participants would’ve preferred another condition. Then, we asked participants to recall a specific day of interaction in terms of how the tool affected their behaviour for that day and for the following day. Finally, participants were asked for their overall thoughts on the tool. The full interview script is shown in Appendix C.
5. Quantitative Analysis
@@
Among the 51 participants in the study, 25 (49%) were from Los Angeles, 10 (19.6%) were from New York, 8 (15.7%) were from Chicago, and 8 (15.7%) were from Houston; 4 (7.8%) were aged 18–24, 15 (29.4%) were aged 25–34, 22 (43.13%) were aged 35–44, 8 (15.7%) were aged 45–54, and 2 (3.9%) were aged over 55; 34 (66.7%) were male, 15 (29.4%) were female, and 1 (2%) was non-binary; 5 (9.8%) had a graduate degree, another 16 (31.4%) had an undergraduate degree, another 11 (21.6%) had a professional degree, and another 19 (37.3%) had a high school degree.
Out of these 51 participants, 44 completed at least one day of interaction with the tool, and 34 completed all 7 days of interaction. Starting from Day 0, Day 7 was reached by 6 (46%) of the baseline Condition (B) participants; 10 (71%) of the daily estimate Condition (D) participants; 7 (58%) of the ranged estimate Condition (R) participants; and 11 (92%) of the ranged and hedged Condition (RH) participants. We show full retention statistics in Figure 3.
In the following sections, we first describe the metrics that we used to measure the participants’ trust and reliance (Section 5.1). Then, we analyse our findings from statistical models for these metrics, specifically those relating to longitudinal effects (Section 5.2), and the effects of specific conditions (Section 5.3).
5.1. Metrics of Trust and Reliance
We measured the trust of participants using self-reported measures, following common practice (Kohn et al., 2021). We used two widely-used instruments for self-reported trust: the Human-Computer Trust Questionnaire (HCT) (Madsen and Gregor, 2000) and the Trust in Automation Scale (TiA) (Jian et al., 2000). HCT measures 5 facets of trust using 5 questions each, while TiA measures both trust and distrust with 12 questions. On Day 0 (pre-interaction), we included 5 items, one taken from each of the HCT’s 5 facets of trust, in the Pre-Survey (Section A.2). On Days 1–7 (during interaction), we included 3 items taken from 3 of the HCT’s 5 facets of trust, in the End-of-Day Survey (Section A.3). On Day 7 (post-interaction), we also included 5 items from the TiA in the Post-Survey (Section A.4), with 3 measuring trust and 2 measuring distrust. All of these questions were presented to participants as 5-point Likert-type scales (Chyung et al., 2017, 2018). From each survey, we computed an overall trust score by first inverting items measuring distrust, if any, and then averaging the Likert-scale responses.
We measured the reliance of participants, i.e. the external behavioural expression of trust, using both self-reported measures (End-of-Day Survey, Question 8; Section A.3) and their actual behaviour of discontinuing study participation. Specifically, we computed it as an ordinal variable with three levels: 1, if the participant indicated in the End-of-Day Survey that they intended to rely on the tool more tomorrow; 0, if the participant indicated that they intended to rely on the tool about the same tomorrow; and -1, if the participant indicated that they intended to rely on the tool less tomorrow, or did not complete the next day’s study activities.
In the following sections, we use this notation to describe the statistical models that we fitted:
pre_trust_score: The Day 0 (Pre-Survey) trust score.
trust_score: The current day’s (End-of-Day) trust score.
reliance: The current day’s (End-of-Day) reliance score.
post_trust_score: The Day 7 (Post-Survey) trust score.
day: The day of interaction with the schedule recommendation tool (1–7).
user_id: A randomly-assigned UUID for each participant, used as random effects.
condition: The participant’s schedule page condition.
estimate_accurate: A binary indicator of whether the participant perceived their earnings to be about the same as the tool’s estimate (End-of-Day Survey, Question 4).
5.2. RQ1: Longitudinal Effects
5.2.1. Effects on Trust
Participants reported a moderately high level of trust in the schedule recommendation tool (). To begin, we analysed how participants’ trust in the schedule recommendation tool changed over time. For the 33 retained participants who completed all 7 days of interaction, Figure 4 (left) shows an upward trend in the mean trust score. To address Research Question 1, we then grouped each day’s trust scores based on whether or not participants perceived the tool’s estimates as being accurate. Figure 4 (right) shows that, on Day 1, perceived accuracy was positively correlated with trust; the interquartile ranges of the trust scores did not overlap between the two groups. This effect was less clear for Days 2–7, where trust scores for the two groups overlapped more extensively.
To further explore the longitudinal effects of perceived accuracy on trust, we fitted a linear mixed model (LMM) for trust_score using the R packages lme4 1.1-35.5 (Bates et al., 2015) and lmerTest 3.1-3 (Kuznetsova et al., 2017). In this model, these longitudinal effects were modelled by the inclusion of the day, perceived accuracy (estimate_accurate), and their interaction as independent variables. We also included the pre_trust_score to adjust for participants’ baseline level of trust in the tool (not on the same scale), and participant IDs as random effects to account for individual variance.
Our model (Table 1 in Section D.2) found that pre-interaction trust was significantly and positively correlated with their daily trust (pre_trust_score: ). Therefore, participants’ baseline trust persisted throughout the their interactions with the tool. Consistent with Figure 4, trust also increased significantly with each passing day (day: ). Also consistent with Figure 4, perceived accuracy was significantly and positively correlated with trust (estimate_accurate: ), but it had less of an impact on trust with each passing day of the user study (day:estimate_accurate: ). This suggests that, by the end of the user study, participants’ trust was based less explicitly on perceived accuracy.
5.2.2. Effects on Reliance
Most participants indicated their desire to maintain their level of reliance on the schedule recommendation tool, corresponding to a reliance score of 0 (). Trust and reliance were not strongly correlated (); some participants consistently expressed high reliance but also lower trust. The mean reliance score appeared to decrease over time, with the mean being lowest on Day 4, but we could discern no clear dependence on perceived accuracy (Figure 6 in Appendix D). To clarify the nature of these longitudinal effects, we fitted another LMM using lme4 and lmerTest. This model was similar to the model for trust, except the reliance score was the dependent variable, and we included the trust_score as an independent variable:
Our model (Table 2 in Section D.2) did not find any significant effects for perceived accuracy (estimate_accurate) or pre-interaction trust (pre_trust_score). However, we did find two significant effects: a negative effect from the day, supporting our initial observation (), and a positive effect from the day:trust_score interaction (). The latter suggests that reliance depended on perceived accuracy indirectly through the trust score. Participants who trusted the tool more were more likely to continue relying on it, with this effect strengthening over repeated interactions even as overall reliance weakened.
5.3. RQ2: Effects of Conditions
5.3.1. Pre-Interaction Trust
Next, we analysed the effects of our design conditions on trust and reliance, starting with pre-interaction trust. We found that the mean pre-interaction trust scores for Conditions (B)/(D)/(R)/(RH) were 3.338, 3.629, 4.183, and 3.367; Condition (B) had the lowest trust, and Condition (R) had the highest.
To verify these observations, we fitted an ordinary least squares (OLS) model for pre-interaction trust using the Python package statsmodels 0.14.2 (Seabold and Perktold, 2010), with the condition as an independent variable. Relative to Condition (B), the daily estimate Condition (D) did not significantly differ in pre-interaction trust (contrast (D)-(B): ); the ranged and hedged estimate Condition (RH) (contrast (RH)-(B): ) did not either. Yet, Condition (R) had significantly higher pre-interaction trust relative to Condition (B) (contrast (R)-(B): ) and Condition (RH) (contrast (RH)-(R): ). Therefore, exposing uncertainty through ranged estimates initially improved participants’ trust. In Section D.3, we report on a similar analysis for post-interaction trust.
5.3.2. Longitudinal Trust and Reliance
Lastly, we assessed the longitudinal effects of the schedule page design condition on trust and reliance. In Figure 5, we show the mean trust and reliance scores of participants in each of the four conditions. The means of the conditions were in most cases similar to the overall mean, with two exceptions: (1) the mean trust and reliance scores for Condition (R) were the highest of all conditions and showed a generally increasing trend; and (2) the trust scores for Condition (D) were lower on Days 5 and 6. To validate these trends, we added the condition and its interaction with the day as independent variables to the LMMs that we fitted in Section 5.2:
Condition (D) decomposed estimated earnings on a daily basis, thus providing information irrelevant to uncertainty. Our models (Table 1 and Table 2 in Section D.2) indicate that this did not significantly improve trust (condition=(D): ) or reliance (condition=(D): ) over the baseline Condition (B). While our trust model did find a significant, positive longitudinal effect for Condition (D) (), Figure 5 suggests that this effect does not represent a practically significant trend.
Condition (R) displayed uncertainty in predicted earnings using ranges of pessimistic and optimistic earnings. Again, our models did not find significant marginal effects for Condition (R) over Condition (B) in trust (condition=(R): ) or reliance (condition=(R): ). However, Condition (R) had nearly significant longitudinal effects for trust (day:condition=(R): ) and also reliance (day:condition=(R): ). This aligns with our observations based on the mean scores in Figure 5 as well as our findings for pre-interaction trust (Section 5.3.1). Therefore, despite exposing uncertainty in the tool’s earnings estimates, the ranges of Condition (R) improved participants’ initial trust and then led them to maintain their trust and reliance over daily interactions.
Condition (RH) added lexical hedging to the ranged earnings estimates in Condition (R). This condition was not significantly different from Condition (B) in terms of marginal or longitudinal effects on trust and reliance. On Day 6, participants in Condition (RH) reported significantly lower reliance than participants in Condition (R) (Figure 6; , ), the only such significant pairwise difference on a daily basis (see Table 3 in Appendix D). Combined with the pre-interaction trust of Condition (RH) being significantly lower than Condition (R) (Section 5.3.1), we conclude that the addition of lexical hedging in Condition (RH) reversed the gains in trust and reliance from Condition (R)’s range-based uncertainty.
6. Qualitative Analysis
Overall, 7 participants completed the exit interview after the longitudinal user study. Three of these were from Condition (D), one was from Condition (R), and three were from Condition (RH):